ABC is a triangle and P, Q are the midpoints of AB, AC respectively. If AB = 2x




and AC = 2y, express the vectors (i) BC, (ii) PQ, (iii) PC, (iv) BQ in terms of x




and y. What can you deduce about the directed line-segments BC and PQ?





1
Expert's answer
2022-02-23T09:32:39-0500

Solution.

"So, \space as \space 2x + BC = 2y, \newline 
BC = 2y - 2x. \newline
As \space x + PQ = y, \newline
PQ = y - x. \newline
As \space x + PC = 2y, \newline
PC = 2y - x. \newline
As \space 2x + BQ = y, \newline
BQ = y - 2x."

Something we can deduce about the directed line-segments "BC" and "PQ" is that

"BC=2PQ" , so they are parallel.

Answer:

"BC=2y−2x, \space
PQ = y - x, \newline
PC = 2y - x, \space
BQ = y - 2x. \newline
BC \space and \space PQ \space are \space parallel."